A study of the effect of massage on boxing performance measured a boxers blood lactate concentration (in mM) and perceived rFind T Statistic and P-value
Answer #1

Answer:

from the given data

Obs. XX YY Xi^2​ Yi^2​ Xi.Yi
1 3.8 8 14.44 64 30.4
2 4.3 8 18.49 64 34.4
3 4.8 11 23.04 121 52.8
4 4.1 12 16.81 144 49.2
5 5.0 12 25 144 60
6 5.3 12 28.09 144 63.6
7 4.2 13 17.64 169 54.6
8 2.4 16 5.76 256 38.4
9 3.9 17 15.21 289 66.3
10 5.3 17 28.09 289 90.1
11 5.8 18 33.64 324 104.4
12 6.0 18 36 324 108
13 5.9 23 34.81 529 135.7
14 6.3 23 39.69 529 144.9
15 5.5 22 30.25 484 121
16 6.5 25 42.25 625 162.5
Sum = 79.1 255 409.21 4499 1316.3

The sum of squares obtained from the table above are:

SSxx = = 400.21 (79.12 18.1594 $$y = = 4499 ) To (255) 434.9375 = 1316.3 -T (70.1) (255) = 55.6438

The slope and y-intercept coefficients are computed using the following formulas: B = SSxy SSxx 55.6438 18.1594 3.0642 Bo = Y

SSE=   (SSxx * SSyy - SS²xy)/SSxx = 264.47

std error ,Se =    √(SSE/(n-2)) = 4.346

slope hypothesis test
Ho:   ß1=   0          
H1:   ß1╪   0          
n=   16              
alpha =   0.1              
estimated std error of slope =Se(ß1) = Se/√Sxx = 4.346 /√ 18.16 = 1.0198
                  
t stat = estimated slope/std error =ß1 /Se(ß1) =  3.0642 / 1.0198 = 3.005
                     
Degree of freedom ,df = n-2=   14              
p-value = 0.0095
decison :    p-value<α , reject Ho .